# Cross-modal Associations in Vision and Language Models: Revisiting the bouba-kiki effect

### v1

@@ Appendix D Grad-Cam visualisation and consistency
Figure 10: Visualising the attention pattern for the text prompt: "The label for this image is <label>" in the both models.
@@ D.1 Prediction consistency
To test whether the models in the second experiment (section 5) change their predictions resulting from different shape positions, we report the ratio of consistency in 3. In this case, we are not interested in the question of whether the models are making a prediction that aligns with a bouba-kiki effect, but are simply interested in the consistency of predictions. Both models are rather consistent in the mappings they make between labels and shapes.
@@
Table 3: The ratio of consistently attending to the same shape in a different position when presented with the same label. Values are averages across image pairs, prompts, and labels. A high ratio means that the corresponding model consistently focuses on the same shape, even when the location is in a different position.

### v2

@@ Appendix D Grad-Cam visualisations, consistency, and additional results
Figure 10: Visualising the attention pattern for the text prompt: ‘The label for this image is <label>’ in both models.
@@ D.1 Prediction consistency
To test whether the models in the second experiment (section 5) change their predictions resulting from different shape positions, we report the ratio of consistency in Table 4. In this case, we are not interested in whether the models make predictions that align with the bouba-kiki effect, but rather in the consistency of their predictions. Both models are rather consistent in the mappings they make between labels and shapes.
@@
Table 4: The ratio of consistently attending to the same shape in a different position when presented with the same label. Values are averages across image pairs, prompts, and labels. A high ratio indicates that the corresponding model consistently focuses on the same shape, even when the target location is different.
D.2 Quantifying model preference
The results presented in section 5 utilise the sum of attention values to quantify the models’ preferences, as this aligns with humans’ holistic perception (Taubert et al.,, 2011; Zhao et al.,, 2016; Wong et al.,, 2011). Though artificial models may show their preference differently. To this end, we additionally experimented with the entropy and centroid-of-attention as quantifiers of model preference.
Comparing the predictions (averaged over all images, labels, and prompts) quantified by the centroid of attention with the sum of attention, we find that the predictions using the centroid of attention strongly overlap with those resulting from the sum of attention (ViT: $85.6\%$ and ResNet: $88.0\%$). Given this considerable overlap, it is not surprising that the results remain highly similar when using this alternative method (Table 5. Yet, for the entropy of attention, the predictions (where the image half with the lower entropy acts as the model’s preference) for ViT overlap strongly ($80.2\%$) but not for ResNet ($21.1\%$). This entropy method, despite yielding different predictions, still results in a slightly lower number of correctly predicted images, with only $48.8\%$ of the shapes correctly identified, compared with $51.9\%$ for the quantification method that uses the sum of attention. So if we were to use entropy-based predictions, the performance would be even worse than predictions based on the sum of attention.
Model Sum Entropy Centroid
ResNet .519 .488 .515
ViT .522 .514 .515
Table 5: The correctness of all model prediction types using different quantifications averaged across all images, labels, and prompts.